# The BPgWSP test: a Bayesian Weibull Shape Parameter signal detection test for adverse drug reactions

### v1

@@ 3.3 Specification of the HDI+ROPE test
The HDI-ROPE test approach is specified with respect to the
@@
We specify the ROPE for each shape parameter as equal-tailed interval (ETI)
@@
based on the quantiles of the shape parameters’ prior distributions under the null-hypothesis at credibility level $\alpha$.
We specify the posterior CI for each shape parameter either as an ETI calculated from the empirical quantiles of the shape parameters’ posterior samples at credibility level $\alpha$ or as an HDI at credibility level $\alpha$ calculated from the posterior sample. An HDI at level $\alpha$ is an interval
[…]
that is with boundaries such that all values inside the HDI are have higher density than values outside the HDI and simultaneously encompasses $(1-\alpha)\cdot 100\%$ of the probability mass (Kruschke,, 2015, p.87).
To control the width of the ROPE as well as the posterior CIs, we apply credibility levels $\alpha$ ranging from $0.5,0.55,...0.95$.
Three different combination rules of the single shape parameters’ test outcomes are defined in Table 3. Option 1 implements the hypothesis formulated in Equation 3. A signal is raised if at least one shape parameter significantly differs from one indicating a non-constant hazard function. Option 2 raises a signal in fewer cases, namely if at least one shape parameter’s ROPE test is positive and the other undecided. Option 3 raises a signal if and only if both single shape parameter tests are positive.
The simulation study is implemented in the statistical software environment R (Team,, 2022) including the packages rstan (Stan Development Team,, 2024) for the Bayesian model setup and posterior sampling, HDInterval (Meredith and Kruschke,, 2022) for calculation of posterior CIs, and BWSPsignal (available on https://github.com/julia-dyck/BWSPsignal) containing functions specifically written for the BPgWSP test development and simulation setup.
@@
Table 3: Three options for the combination of the single test outcomes for shape parameter $\nu$ and powershape parameter $\gamma$. Option 1 raises a signal if at least one parameter significantly differs from one. Option 2 raises a signal in fewer cases. Option 3 raises a signal only if both single parameter tests are positive.
@@ 3.4 Evaluation criteria for the simulation results
The evaluation of the simulation study is performed in three steps. The first step consists of selecting a combination of prior distributions (see Table 2) that provides reliable posterior estimates. We will reject all combinations of prior distributions that lead to cases of non-convergence in the posterior estimation or that tend to higher execution times or small effective posterior sample sizes. According to Kruschke, (2015), a posterior sample is recommended to be used for HDI+ROPE testing if the effective sample size is at least $10000$.
The second step includes a classification of tuning parameter combinations for the HDI+ROPE test and combination rule based on a sensitivity and specificity driven performance measure. The classification is restricted to prior distributions selected in step one. The performance of the implemented test setups (depending on ROPE and posterior CI definition, credibility level and combination rule) averaged over varying simulation scenarios (sample size of data, background rate, ADR rate, expected event-time) are measured in terms of the area under the curve (AUC) of the receiver operating characteristic graph with one threshold (Fawcett,, 2004; Lloyd,, 1998) based on equal numbers of ADR and no-ADR scenarios (and thereby equal numbers of true positives and negatives). The AUC is calculated using the R package ROCR (Sing et al.,, 2005).
In the third step, the BPgWSP test with the tuning parameters chosen in step two is investigated further. The influence of the sample scenario on the AUC performance is presented in terms of the AUC stratified with respect to sample size, ADR rate, and expected event-time. Finally, the robustness to miss-specification is explored by examining the AUC when the prior belief differs from the truth because the expected event-time is incorrectly specified.
@@ 3.5 Results of the simulation study
Per prior distributional choice (see Table 2) 8400 posterior estimations (resulting from 336 simulation scenarios divided by number of distribution choices times 100 repetitions) were conducted. In the following, the prior distributions are compared with respect to convergence issues, execution times and effective sample sizes.
Non-convergence in posterior estimation:
Only few posterior estimations failed due to non-convergence of the MCMC. This occurred for $3.6\%$ of the posterior estimations given a fix-gam-gam prior and for $10.2\%$ of the posterior estimations when a gam-gam-gam prior was used. No implications for convergence issues (in form of an error message and canceled execution of posterior estimation) came up when a fix-log-log or log-log-log prior was specified. As they seem to provide smooth estimation without the need to adjust the MCMC sampler, we suggest to use a fix-log-log or a log-log-log prior distribution.
Execution times:
Figure 5: Boxplots of running times split with respect to prior distributional choice. The graph is cut on the y-axis at 210 minutes (3h) although the boxplot for the gam-gam-gam prior reaches as maximum at about 45 hours.
A boxplot of the execution times for posterior estimation stratified with respect to the prior distribution is shown in Figure 5. All distribution choices lead to acceptable median execution times of 8 min for fix-log-log models, 13 min for log-log-log models, 10 min for fix-gam-gam models, and 40 min for gam-gam-gam models. While the maximal running times that occurred during the simulation stayed under about 1 h given a fix-log-log, 2 h given a fix-gam-gam, or 3 h given a log-log-log distribution, the gam-gam-gam prior produced execution times of up to 45 hours.
Effective sample sizes:
@@
Figure 6: Boxplots of estimated effective sample size split with respect to prior distributional choice for parameter $\nu$ and $\gamma$.
In $73\%$ of the posterior estimations for $\nu$ and $55\%$ of the posterior estimations for $\gamma$ does the log-log-log model generate a posterior sample with an effective sample size $>10000$. The fix-log-log, fix-gam-gam and gam-gam-gam prior distributions provide only $\approx 25-28\%$ simulations for $\nu$ and $21-26\%$ for $\gamma$ fulfilling the rule of thumb (see Figure 6). This makes the log-log-log prior preferable compared to the alternative prior distributions.
The log-log-log distribution provides the best results in terms of convergence of the posterior estimation, adequate running times and sufficiently high effective posterior sample size in most cases. Consequently, we now proceed to evaluate the simulation results of the the log-log-log prior distribution.
Ranking of test specifications based on the AUC:
We rank all combinations of ROPE and posterior CI specifications, credibility levels and combination rules based on the AUC averaged over all sample scenarios given a correct prior belief. Table 4 shows the top ten test specifications in terms of AUC. The highest AUC value of $0.89$ is achieved with an $80\%$ Lognormal ETI as ROPE, an $80\%$ posterior HDI as posterior CI and option 2 for the combination rule (see Table 3).
rank ROPE posterior CI combination rule AUC
1 80% ETI 80% HDI option 2 0.890
2 80% ETI 80% ETI option 2 0.873
3 80% ETI 80% ETI option 3 0.838
4 80% ETI 80% HDI option 3 0.834
5 75% ETI 75% HDI option 3 0.823
6 75% ETI 75% ETI option 3 0.802
7 85% ETI 85% HDI option 2 0.693
8 85% ETI 85% ETI option 2 0.691
9 90% ETI 90% HDI option 1 0.653
10 85% ETI 85% HDI option 1 0.643
@@
Table 4: Ranking of top ten BPgWSP test specifications according to AUC averaged over all sample scenarios given a correct prior belief.
Effect of simulation scenario on AUC performance:
@@
The influence of the sample size, the proportion of ADR events and the expected event-time of the ADR (with correctly specified prior belief) is summarized in Table 5. The BPgWSP test works better with increasing sample size reaching an AUC of $98.9\%$ given a sample size of 5000. An increase in the proportion of ADR events seems to improve detectability of the ADR very slightly. The test performance is higher for ADRs occurring around the first or third quarter of the observation period with an AUC of $94\%$, but still provides an AUC of $78\%$ for ADRs occurring in the middle of the observation period.
@@
1st quarter of observation period 0.946
2nd quarter of observation period 0.782
3rd quarter of observation period 0.942
Table 5: Average AUC performance of the BPgWSP test (prior distribution: log-log-log, ROPE: $80\%$ Lognormal confidence interval, credibility interval: $80\%$ highest posterior density region, combination rule option 2) grouped by sample size, proportion of ADR events, and expected time of ADR events.
Robustness to miss-specification:
If the specification of the expected event-time is off by one quarter of the observation time, the AUC reduces to $78\%$ on average. If the a priori expected event-time is off by two quarters, the AUC performance drops to $47\%$. An AUC performance of $79\%$ is provided in case no ADR is suspected a priori.
For application of the BPgWSP test to a range of drug-AE pairs (in an automated manner), we recommend to set up the tuning parameters as follows: an $80\%$ Lognormal ETI for the ROPE, an $80\%$ posterior HDI for the posterior CI and option 2 for the combination rule. In the next section, the developed method will be applied to real data for demonstration purposes using the recommended tuning for the BPgWSP test.
@@
correct specification 0.890
one quarter off 0.780
two quarters off 0.469
no ADR assumed 0.788
@@ 4 Case study
The purpose of this case study is to illustrate the BPgWSP test derived in this work with a real dataset. Conclusions from this case study should not be used to inform clinical practice. The dataset was originally created from the Health Improvement Network (THIN) database (THIN,, 2024) to assess the association between the prescription of bisphosphonates and the occurrence of Carpal tunnel syndrome in menopausal Women (Carvajal et al.,, 2016). The THIN database contains primary care data from patients in the UK.
The BPgWSP test is applied to investigate the association between bisphosphonates and four AEs occurring in the population. In addition to the exposed cohort, the dataset contains patients not exposed to bisphosphonates (control group). A control group is not required to apply the BPgWSP test. Nevertheless, it provides us with the opportunity to explore the type one error, i.e. the test results for the control group.
The data preparation has been described in Sauzet et al., (2013). To facilitate the description of the data, we recall the descriptive statistics provided in Sauzet et al., (2013). All steps are conducted for the exposed and the control groups. We present the sample size and number of events as well as the mean event-time (and standard deviation). A histogram of event-times gives insights into the distribution over time. A smoothed non-parametric estimate of the hazard function provides information on the trend of risk for an AE to occur over time.
The BPgWSP test is applied with the recommended tuning parameter resulting from the simulation study (log-log-log priors, 80% Lognormal confidence interval for the ROPE, 80% HDI for the posterior CI, option 2 for the combination rule). Selectively, we compare our test outcomes to the test results obtained with the non-Bayesian approaches illustrated in Sauzet et al., (2013) and Sauzet and Cornelius, (2022).
@@ 4.1 Headache
Figure 7: Graphics for exploratory analysis of the occurrence of headache over 21 days. The graphic derived is based on Sauzet et al., (2013, Fig. 1).
@@
The data is censored after 21 days. In this time, 21 among 19791 patients $(0.11\%)$ exposed to bisphosponates experienced headaches. The day of occurrence is on average $14\;(5.9)$. The histogram of event-times (Figure 7 a.1) shows that the number of events is slightly higher in the second half of the observation period. The non-parametric smooth hazard estimate (Figure 7 b.1) indicates a more pronounced increase of risk in the last seven days.
@@
prior belief about prior means test outcome test outcome
@@
none $(1,1,1)$ signal -
1st quarter of OP (5.25 days) $(2.5,0.5,1)$ signal -
2nd quarter of OP (10.5 days) $(3,3,5)$ signal -
3rd quarter of OP (15.75 days) $(18,5,1)$ signal -
Table 7: Test outcomes under various prior beliefs for headache based on exposed and control group.
Figure 8: Prior distribution and posterior sample histogram obtained from Bayesian model under the prior belief that headache is no ADR. The test results in row one of Table 7 are concluded from the alignment of the ROPE and HDIs and combination according to combination rule 2.
@@
The prior specifications and results of the BPgWSP test are shown in Table 7. The signal detection tool is applied to the data censored at 21 days under the prior beliefs that the symptom occurrence is expected at the first, second or third quarter of the observation period, as well as a prior belief of no association. Independent of the prior specification including the assumption of no association, a signal is consistently raised for the exposed group and no signal for the control group. This is in line with the results of the non-Bayesian approach.
The intermediate shape parameters’ HDI+ROPE test outcomes are illustrated in Figure 8 for the exposed and control group under the prior belief of no association. In the case of the exposed cohort, both single parameter tests reject the null-hypothesis as the HDI does not lie within the ROPE) and combination scheme option 2 concludes that a signal is raised. In contrast, the control group’s HDI+ROPE test for $\gamma$ returns a rejection of the null-hypothesis while the HDI+ROPE test for $\nu$ returns an acceptance. According to the combination scheme option 2, no signal is raised.
@@ 4.2 Musculoskeletal pain
Figure 9: Graphics for exploratory analysis of the occurrence of musculoskeletal pain over one year. The graphic derived is based on Sauzet et al., (2013, Fig. 2).
@@
The data is censored after one year in which 207 of 19777 ($1\%$) exposed patients reported musculoskeletal pain. The average event-time to occurrence is 160 days $(99)$. The histogram of event-times (Figure 9 a.1) shows a slightly higher density in the first half of the observation period. The smoothed hazard estimate (Figure 9 b.1) indicates a slight increase in risk during the first four months followed by a decrease.
In the control group musculoskeletal pain was reported 207 times from 39634 patients ($0.5\%$). According to the histogram (Figure 9 a.2) slightly more cases are reported in the first and third month but evenly distributed for the rest of the time. The hazard estimate (Figure 9 b.2) looks almost constant.
prior belief about prior means test outcome test outcome
@@
none $(1,1,1)$ signal -
1st quarter of OP (91 days) $(1,0.207,1)$ signal -
2nd quarter of OP (183 days) $(20,5.5,14)$ signal -
3rd quarter of OP (274 days) $(300,4,1)$
Table 8: Test outcomes under various prior beliefs for musculoskeletal pain based on exposed and control group.
@@
The results of the BPgWSP test applied to the data censored at 365 days under the prior beliefs of symptom occurrence at the first, second or third quarter of the observation period or of no association are shown in Table 8. A signal is raised for the exposed group under all priors except a prior belief around the third quarter (day 274). For the control group, no signal is raised regardless of the prior choice.
In the case of musculoskeletal pain, the test results depend on the prior specification. Here we have a long observation period and a small number of cases which potentially decreases the performance of the BPgWSP test. Given that most cases occur at the beginning of the observation period the prior specification for the last observation quarter does not lead to a signal. This presumably occurs because data and prior belief provide contrasting information.

### v3

@@ 3.3 Specification of the HDI+ROPE test
The Bayesian HDI+ROPE test approach is specified with respect to the
@@
We specify the ROPE for each shape parameter $\nu$ and $\gamma$ as equal-tailed interval (ETI)
@@
based on the quantiles $q$ of the shape parameters’ prior distributions under the null-hypothesis at credibility level $\alpha$.
The posterior CI at credibility level $\alpha$ for each shape parameter is set to be an ETI (see Equation 9) calculated from the empirical quantiles of the shape parameters’ posterior samples or an HDI estimated based on the posterior sample. An HDI at level $\alpha$ is an interval
$HDI(\nu)=\{\nu\;|\;p_{1}(\nu)\geq w\}\\ \text{ with }w\in[0,1]\text{ such that}\int_{\nu\;|\;p_{1}(\nu)\geq w}p_{1}(\nu|t)\;d\nu=1-\alpha,$ (10)
that is with boundaries such that all values inside the HDI have a higher density than values outside the HDI and simultaneously encompasses $(1-\alpha)\cdot 100\%$ of the probability mass.19
To control the width of the ROPE as well as the posterior CIs, we apply credibility levels $\alpha$ ranging from 0.5,0.55,0.950.5,0.55,...0.95.
@@
Table 2: Three options for the combination of the single test outcomes for shape parameter $\nu$ and powershape parameter $\gamma$. Option 1 raises a signal if at least one parameter significantly differs from one. Option 2 raises a signal in fewer cases, namely when one parameter test rejects the null hypothesis $H_{0}$ and the other rejects $H_{0}$ or is undecided. Option 3 raises a signal only if both single parameter tests reject $H_{0}$.
Three different combination rules of the single shape parameters’ test outcomes are defined in Table 2. Option 1 implements the hypothesis formulated in Equation 3. A signal is raised if at least one shape parameter significantly differs from one indicating a non-constant hazard function. Option 2 raises a signal in fewer cases, namely if the null-hypothesis is rejected for at least one shape parameter and rejected or undecided for the other. Option 3 raises a signal if and only if both single shape parameter tests reject the null-hypothesis.
The pWSP test is constructed according to Sauzet and Cornelius.6 Based on the maximum likelihood estimate (MLE) the confidence interval at confidence level $\alpha$
$[\hat{\nu}_{MLE}\pm z_{(1+\alpha)/2}\cdot sd(\hat{\nu}_{MLE})]$ (11)
is calculated and used for significance testing for each shape parameter. We apply confidence levels 0.5,0.55,0.90.5,0.55,...0.9 and 0.91,0.990.91,...0.99 as well as 0.991,0.9990.991,...0.999 to take into account the relevant levels according to Sauzet et al. 22 When shape parameters cannot be estimated (e.g. due to convergence issues), this is considered as no rejection of the single parameter null-hypothesis.
A signal is raised if both parameter estimates differ significantly from the null value.
The simulation study is implemented in the statistical software environment R 23 including the packages rstan 24 for the Bayesian model setup and posterior sampling, HDInterval 25 for calculation of posterior CIs, and BWSPsignal 26 containing functions for the BPgWSP test and simulation framework.
@@ 3.4 Evaluation criteria for the simulation results
The evaluation of the simulation study is performed in three steps. First, prior distributions that show convergence issues, extremely high execution times, or small posterior effective sample sizes (ESS) are excluded. The model is considered to have not converged whenever the rstan::sampling function returns an error. We consider the execution time extreme if the maximum is 10 hours or more per model. Finally, following Kruschke’s recommendation for HDI+ROPE testing, we consider prior distributions exceeding a posterior ESS of $10,000$ or more on average as sufficient for testing. 19, 21
Second, the tuning parameter combinations are classified based on the test sensitivity and specificity under correct prior beliefs. The performance of the implemented test setups averaged over varying simulation scenarios is measured in terms of the area under the curve (AUC) of the receiver operating characteristic (ROC) graph with one threshold 27, 28 based on equal numbers of ADR and no-ADR scenarios. In addition to the AUC value, the ROC curve allows further investigation of the top tests by displaying the AUC depending on the false and true positive (FP, TP) rates. The AUC is calculated using the R package ROCR. 29
Third, to explore the effect of sample size, ADR rate, and expected event-time on the test performance, we present grouped AUCs of the best test. In case of a BPgWSP test, the robustness to miss-specification is explored by examining the AUC when the prior belief differs from the truth and the expected event-time is therefore incorrectly specified.
@@ 3.5 Results of the simulation study
We compare the prior distributions with respect to convergence issues, execution times, and ESSs. There were no convergence issues.
Figure 5: Boxplot comparing execution times in minutes for two prior distributions, gamma (pgw-$g$) and lognormal (pgw-$l$), with lognormal yielding higher maximum run times of up to 3 h, but both prior distributions producing similar median times of 16 Minutes with gamma and 17 Minutes with lognormal prior.
@@
Figure 6: Boxplots of estimated effective sample sizes (n_eff) split with respect to prior distributional choices gamma (pgw-$g$) and lognormal (pgw-$l$) for parameter $\nu$ (shape1 in red) and $\gamma$ (shape2 in blue). The recommended threshold for the effective sample size is marked with a horizontal dashed line at 10 00010\,000.
A boxplot of the execution times for posterior estimation stratified with respect to the prior distribution is shown in Figure 6. The models produce median execution times of 16 Minutes with gamma priors, and 17 Minutes with lognormal priors. The maximal running times stayed under 1 h 40 min given a gamma prior and under 3h given a lognormal prior. Gamma priors lead to ESSs of >10 000>10\,000 in $85\%$ of the posterior estimations for $\nu$, and $68\%$ for $\gamma$. For lognormal priors these number are $73\%$ for $\nu$ and $55\%$ for $\gamma$ (see Figure 6).
Based on execution time and ESS results, both prior distribution choices are acceptable for the modeling part of the BPgWSP.
Figure 7: Roc curves representing the top one (left) and top ten (right) BPgWSP tests’ performances.
rank prior ROPE posterior CI combination AUC FP rate TP rate
distribution rule
1 lognormal 80% ETI 80% HDI option 2 0.900 0.173 0.974
2 lognormal 80% ETI 80% ETI option 2 0.876 0.226 0.977
3 lognormal 80% ETI 80% ETI option 3 0.844 0.006 0.694
4 lognormal 80% ETI 80% HDI option 3 0.837 0.001 0.675
5 lognormal 75% ETI 75% HDI option 3 0.823 0.326 0.971
6 gamma 93% ETI 93% HDI option 2 0.810 0.356 0.976
7 lognormal 75% ETI 75% ETI option 3 0.808 0.364 0.980
8 gamma 93% ETI 93% ETI option 3 0.804 0.003 0.611
9 gamma 93% ETI 93% ETI option 2 0.803 0.371 0.978
10 gamma 93% ETI 93% HDI option 3 0.797 0.002 0.597
@@
Table 3: Top ten BPgWSP and pWSP test specifications ranked based on AUC. AUC, FP rate and TP rate are averaged over all sample scenarios given a correct prior belief.
Table 3 shows the top ten test specifications in terms of AUC along with the corresponding FP and TP rates. Figure 7 displays the ROC curves of the top one (left) and top ten (right) test specifications. The highest AUC value of $0.9$ is achieved with an $80\%$ lognormal ETI as ROPE, an $80\%$ posterior HDI as posterior CI and option 2 for the combination rule. The corresponding ROC curve exhibits a TP rate of $0.97$ and a FP rate of $0.17$, indicating a trade-off that prioritizes the detection of TPs. The best pWSP test (frequentist) with confidence level of $55\%$ is ranked in 11th place with an AUC of $0.75$ under the given simulation conditions.
@@
The influence of the sample size $N$, the proportion of ADR events and the expected event-time of the ADR on the AUC under a correctly specified prior belief is summarized in Table 4. The BPgWSP test works better with increasing $N$ reaching an AUC of $0.978$ given an $N$ of 5000. An increase in the proportion of ADR events from $5\%$ to $10\%$ of $N$ increases the AUC very slightly from $0.888$ to $0.913$. The test performance is higher for ADRs occurring around the first or third quarter of the OP with an AUC of $0.969$ or $0.946$, respectively, but still provides an AUC of $0.786$ for ADRs occurring in the middle of the OP.
@@
1st quarter of observation period 0.969
2nd quarter of observation period 0.786
3rd quarter of observation period 0.946
Table 4: Average AUC performance of the BPgWSP test with lognormal prior distribution, $80\%$ lognormal confidence interval as ROPE, $80\%$ HDI as posterior CI, and combination rule option 2 grouped by sample size, proportion of ADR events, and expected time of ADR events given a correct prior belief.
If the specification of the expected event-time is off by one quarter of the observation time, the AUC reduces to $0.783$ on average. If the a priori expected event-time is off by two quarters, the AUC performance drops to $0.489$. An AUC performance of $0.787$ is provided in case no ADR is suspected a priori (see Table 5).
@@
correct specification 0.900
one quarter off 0.783
two quarters off 0.489
no ADR assumed 0.787
@@
For application, we recommend setting the tuning parameters as follows: an $80\%$ lognormal ETI for the ROPE, an $80\%$ posterior HDI for the posterior CI and option 2 for the combination rule.
@@ 4 Case study
The purpose of this case study is to illustrate the BPgWSP test with a real dataset. Conclusions from this section should not be used to inform clinical practice. The dataset was originally created from the Health Improvement Network (THIN) database 30 to assess the association between the prescription of bisphosphonates and the occurrence of carpal tunnel syndrome in menopausal women.31 The THIN database contains primary care data from patients in the UK.
The BPgWSP test is applied to investigate the association between bisphosphonates and four AEs occurring in the population. In addition to the exposed cohort, the dataset contains patients not exposed to bisphosphonates (control group). A control group is not required to apply the BPgWSP test, but allows to explore potential type I errors. The data is described in more detail in Sauzet et al. 32
We illustrate the BPgWSP test using four different prior beliefs for the lognormal parameters reflecting an a priori expected ADR event-time at the first, second, or third quartile of the OP, or none. Given the lack of prior knowledge, we choose priors based on visualizing the corresponding hazard. Otherwise, we used the same model set-up as in the simulation study.
If the ESS is lower than 10 00010\,000 or the scale reduction factor21 $\hat{R}$ far from one, a larger posterior sample is generated. The BPgWSP test is applied with the recommended tuning of the simulation study.
The R code and results are stored on https://osf.io/p8hvr/files/osfstorage. See the supplementary material for a more detailed description of the analysis.
@@ 4.1 Headache
Figure 8: Graphics for exploratory analysis of the occurrence of headache over 21 days. The graphic derived is based on Sauzet et al., Fig. 1. 32
@@
The data is censored after 21 days. In this time, 21 among 19791 patients $(0.11\%)$ exposed to bisphosponates experienced headaches. The day of occurrence is on average 14(sd=5.9)14\;(sd=5.9). The histogram of event-times (Figure 8 a.1) shows that the number of events is slightly higher in the second half of the OP. The non-parametric smooth hazard estimate (Figure 8 b.1) indicates a more pronounced increase of risk in the last seven days.
@@
prior belief about a priori $E[t]$ prior means test outcome test outcome
@@
none - $(1,1,1)$ - -
1st quarter of OP 5.25 days $(2.5,0.5,1)$ - -
2nd quarter of OP 10.5 days $(3,3,5)$ signal -
3rd quarter of OP 15.75 days $(18,5,1)$ signal -
Table 6: Test outcomes under various prior beliefs for headache based on exposed and control group. Prior standard deviation is set to 10 for all parameters in all prior belief cases.
@@
A signal is raised for the exposed group under the prior belief that ADR occurs towards the middle or end of the OP. No signal is raised under the prior of no ADR or ADR at the beginning of the OP (see Table 6). For the control group, no signal is raised independent of the prior specification.
Headache is a known ADR of bisphosphonates. 33 The BPgWSP test detects a signal under the prior belief of ADR occurrence towards the middle or end of the OP. These beliefs presumably amplify the signal in the data already indicating a higher hazard rate towards the end of the OP (see Figure 8 b.1) while the other prior settings might fail to detect a signal due to the contradictory prior and data information.
@@ 4.2 Musculoskeletal pain
Figure 9: Graphics for exploratory analysis of the occurrence of musculoskeletal pain over one year. The graphic derived is based on Sauzet et al., Fig. 2. 32
@@
The data is censored after one year in which 344 of 19777 ($1.7\%$) exposed patients reported musculoskeletal pain. The average event-time is 160 days $(sd=99)$. The histogram of event-times (Figure 9 a.1) shows a slightly higher density in the first half of the OP. The smoothed hazard estimate (Figure 9 b.1) indicates a slight increase in risk during the first four months followed by a decrease.
In the control group musculoskeletal pain was reported 366 times from 39634 patients ($0.9\%$). According to the histogram (Figure 9 a.2) slightly more cases are reported in the first and third month and evenly distributed for the rest of the time. The hazard estimate (Figure 9 b.2) looks almost constant.
prior belief about a priori $E[t]$ prior means test outcome test outcome
@@
none - $(1,1,1)$ signal -
1st quarter of OP 91 days $(1,0.207,1)$ signal -
2nd quarter of OP 183 days $(20,5.5,14)$ signal -
3rd quarter of OP 274 days $(300,4,1)$ - -
Table 7: Test outcomes under various prior beliefs for musculoskeletal pain based on exposed and control group. Prior standard deviation is set to 10 for all parameters in all prior belief cases.
@@
A signal is raised for the exposed group under all priors except a prior belief around the third quarter (day 274). For the control group, no signal is raised regardless of the prior choice (see Table 7).
Musculoskeletal pain is a recognized ADR that can occur anytime after the beginning of bisphosphonate intake.34 Here, the test result depends on the prior specification. A long OP and a small number of cases potentially decrease the performance of the BPgWSP test when the prior contrasts sharply with the data, as is the case under the end-prior here. Moreover, the result of no signal for the exposed group that was based on the model that was not rerun might indicate the need for more iterations, even when $\hat{R}$ and ESS suggest convergence.